Find all left cosets of the subgroup {ρ02} = {e,(1,3)(2,4)} in D4



1
Expert's answer
2021-04-15T07:25:19-0400

Remind that "D_4=\{\rho_0=e,\rho_1=(1\ 2\ 3\ 4), \rho_2=\rho_1^2, \rho_3=\rho_1^3, \sigma=(1\ 3), \sigma\rho_1, \sigma\rho_1^2, \sigma\rho_1^3\}" , "\sigma^2=\rho_1^4=(\sigma\rho_1)^2=e". Given an element "g\in D_4", the left coset in "D_4", generated by the element g and the subgroup "H=\{e,\rho_1^2\}", is the set "\{gh:h\in H\}" . Two left cosets gH and g'H are the same, if and only if "g'\in gH" , if and only if "g^{-1}g'\in H" . Let calculate.

"eH=H=\{e,\rho_1^2\}=\rho_1^2H"

"\rho_1 H=\{\rho_1,\rho_1^3\}=\rho_1^3H"

"\sigma H=\{\sigma, \sigma\rho_1^2\}=\sigma\rho_1^2H"

"\sigma\rho_1 H=\{\sigma\rho_1, \sigma\rho_1^3\}=\sigma\rho_1^3H"

The number of different left cosets is 4. This is an index of the subgroup H in "D_4" .